Suppose you have 77$ for Tickets, and you want a total of 7 tickets. You want some weekday bleachers seats and some weekend lawn seats both are $11. How many of each type of ticket can you buy and explain. I also have to write it in systems.
Answer by stanbon(75887)

Suppose you have 77$ for Tickets, and you want a total of 7 tickets. You want some weekday bleachers seats and some weekend lawn seats both are $11. How many of each type of ticket can you buy and explain. I also have to write it in systems.
----
# of tickets you can buy = 77/11 = 7
----
Equation:
B + L = 7
---
Solutions: 0/7 ; 1/6 ; 2/5 ; 3;4 ; 4/3 ; 5/2 ; 6/1 ; 7/0
==============================
Cheers,
Stan H.